OpenStudy (anonymous):

The mean average temperature for Fairbanks, Alaska, is 27° F. The monthly average temperatures vary between 36.5° above and below this value. If t = 0 represents February, the phase shift of the sine function is 2. a. Write a model for the average monthly temperature in Fairbanks, Alaska. b. According to your model, what is the average temperature in November?

OpenStudy (anonymous):

so you're iven the amplitude as 36.5

OpenStudy (anonymous):

the average is say... your starting point

OpenStudy (anonymous):

equation of sine in terms of a,b,c \[sin(ax+b)+c\]

OpenStudy (anonymous):

actually \[asin(bx+c)+d\]

OpenStudy (anonymous):

a is your amplitude, d is your shift upwards and downwards,, so now putting those in you get...

OpenStudy (anonymous):

now with phase shifts , which is c so now youhae \[36.5=a\] \[27=d\] \[2=c\]

OpenStudy (anonymous):

36.5sin(bx+2)+27?

OpenStudy (anonymous):

i believe so...

OpenStudy (anonymous):

Sorry for being so like dumb right now, it;s 4am :l and thank you n.n

OpenStudy (anonymous):

it's 4 here also

OpenStudy (anonymous):

ugh I just looked at the answers and t has to fit in there :l want the answers? And why are you awake? Just wanting to help people? cx

OpenStudy (agent0smith):

Isn't x = t here? I think you're just using x in your formula instead of t...

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

sorry caps lol

OpenStudy (anonymous):

but the period part is what i'm wondering upon

OpenStudy (anonymous):

Oh that would make more sense.

OpenStudy (agent0smith):

Period is one month... i think. Or two months maybe. "average temperature for Fairbanks, Alaska, is 27° F. The monthly average temperatures vary between 36.5° above and below this value." I think this is saying that one month is 36.5 above, the next is 36.5 below... making the period 2 months.

OpenStudy (anonymous):

yeah worded very vague

OpenStudy (anonymous):

but in that case that wouldn't make it 2 months a period... it'd make it 4

OpenStudy (anonymous):

possible answers: y=18.3sin[t/6pi + pi/3] +27 y=27sin[t/6pi + 2pi/3] +36.5 y=18.3sin[t/12pi - pi/6] +27 y=36.5sin[t/6pi - pi/3] +27

OpenStudy (anonymous):

Feb-27, March-27+36 april-27, may-27+36 june 27

OpenStudy (agent0smith):

Yeah i'm really not sure what the period is meant to be... but yeah four months might actually be right

OpenStudy (anonymous):

I'm just confused at this point honestly.

OpenStudy (anonymous):

was the phase shift suposed to be in pi and it didn't show up?

OpenStudy (agent0smith):

Looks like it has to be last option. That's the only one with amplitude 36.5.

OpenStudy (anonymous):

nope, all of it is there :l and yeah I was thinking the same thing @agent0smith

OpenStudy (agent0smith):

y=36.5sin[t/6pi - pi/3] +27 from that one, phase shift = (pi/3)/(pi/6) = 2 d = 27 A = 36.5 It all fits with the question.

OpenStudy (anonymous):

oh I see c: